OpenStudy (anonymous):

Write an expression in simplest form for the area of a circle with a radius of 4r. Use 3.14 for pi. A. 3.24r^2 B. 50.24r^2 C. 12.56r^2 D. 12.56r

OpenStudy (hoblos):

area of a circle= pi*R^2 = 3.14*16r^2 = 50.24r^2